a rancher is planning to put up 220 yards of fencing. IN the morning she puts up 80 yards, and in the afternoon she puts up 40% of the remaining fence. What percent of the fence did she put up that day?
Answer by lefty4ever26(59)

Starts with 220 yards of fencing. In the morning 80 yds is put up.
SO,
220-80=140. Then 40% of remaining (140) is put up.
140*.40=56. What is percentage put up in that one day??
140+56=196 To find percentage divide amount used by beginning amount.
196/220=.89 or 89%
Good Luck!